Given the following, calculate expiratory reserve volume TV = 450,000 microL VC = 6.2 L IRV...

  1. Given the following, calculate expiratory reserve volume

TV = 450,000 microL

VC = 6.2 L

IRV = 1700 mL

  1. How does vital capacity normally vary with age?
  1. How does residual volume normally vary with age?
  1. How does residual volume and vital capacity vary with height?
  1. During an asthma attack, intense bronchoconstriction can occur. How would this affect expiratory reserve volume?
  1. What respiratory muscle is impaired by pregnancy? Explain how.
  1. In obstructive lung diseases expiration is difficult.
  1. How would this affect residual volume?
  1. How would this affect vital capacity?
  1. Mr. Smith’s tidal volume is 1/8 of his vital capacity. His expiratory reserve volume is 1/2 of his inspiratory volume. His residual volume is 1/4 of his lung capacity. If his total lung capacity is 6400 cc, solve for:
  1. Tidal volume
  2. Expiratory reserve volume
  3. Inspiratory reserve volume
  4. Vital capacity
  5. Residual volume
  1. Name the subdivisions of the larynx from the most superior to the most inferior respectively.
  1. Describe the epithelium superior to the vocal folds and inferior to the vocal folds?
Answer #1

The answer to the first question is as follows

ERV = exiratory reserve volume

The formula to calculate VC is

VC= TV+IRV+ERV

lets substitute the values available in the question, and convert all units to ml for ease of calculation.

VC= 6.2L = 6200ml

TV = 450000 microL = 450ml

IRV = 1700ml

on substituting the values,

6200ml = 450ml+1700ml+ERV

6200= 2150+ERV

ERV=6200-2150 = 4050ml

PS: according to policy only first question is solved since the questions to be solved is not mentioned, please mention the questions to be solved next time to help us.